Find the following f[g(x)] if f(x)=3x-7 and g(x)=x^3-2x^2-x+1
Answer by John10(297)

Find the following f[g(x)] if f(x)=3x-7 and g(x)=x^3-2x^2-x+1
---------------------------------
We know that f(x) = 3x - 7
To find (fog)(x)= f(g(x))= f(x^3-2x^2-x+1) = 3(x^3-2x^2-x+1) - 7
= 3x^3 -6x^2 -3x -4
The above step just replace x by g(x).